OpenStudy (anonymous):

Use mathematical induction to prove that ((n^2)+5)n is divisible by 6 for all integers n>=1.

OpenStudy (jamesj):

So, write it out carefully. Let P(n) be the statement: ((n^2)+5)n is divisible by 6 You want to show that Step 1. P(1) is true Step 2. That P(k) ==> P(k+1) for all integers k greater than or equal to 1. Step 3. Then conclude by the Principle of Mathematical Induction that P(n) is true for all n greater than or equal to 1.

OpenStudy (jamesj):

Given those three steps, which one or ones are you having trouble with?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

so for step one I have:

OpenStudy (anonymous):

\[p(1): [((1)^{2}+5)(1)] / 6 MOD 0 ->6/6 MOD 0\]

OpenStudy (anonymous):

Is it ok to use % (MOD) to prove this?

OpenStudy (anonymous):

When I use mod = 0, I just think: remainder of 0

OpenStudy (anonymous):

You here hero?

OpenStudy (jamesj):

no, don't use mod. You're making it too complicated.

OpenStudy (jamesj):

Just look at it. When n = 1, what does ((n^2)+5)n equal?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

so it is divisible for p(1)

OpenStudy (jamesj):

Right and obviously 6 is divisible by 6 so the statement P(1) is true.

OpenStudy (anonymous):

so assume p(n) is divisivle by 6

OpenStudy (anonymous):

Then prove p(n)=>p(n+1)

OpenStudy (jamesj):

Step 2: Now suppose P(k) for k >= 1. I.e., k(k^2 + 5) = 6m for some integer m

OpenStudy (jamesj):

Now you want to show that (k+1)( (k+1)^2 + 5 ) = 6p for some other integer p using what I just wrote down.

OpenStudy (jamesj):

This shows that P(k) ==> P(k+1)

OpenStudy (anonymous):

Right

OpenStudy (jamesj):

So now manipulate (k+1)( (k+1)^2 + 5 )

OpenStudy (anonymous):

So what i've done so far is facotr and foil and what not

OpenStudy (anonymous):

to get:

OpenStudy (anonymous):

\[n ^{3}+2n ^{2}+6n+6\]

OpenStudy (jamesj):

No

OpenStudy (jamesj):

and write it in k.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Did I make an error, or should I not multiply out

OpenStudy (jamesj):

Yes you made an error. But in any case, you want to exploit terms that look like k(k^2 + 5) so completely expanding out might not be the best strategy.

OpenStudy (anonymous):

So if k = 6m , where m is an integer, then k is divisible by 6

OpenStudy (jamesj):

No, that's not what P(k) says. P(k) says that k(k^2 + 5) is divisible by 6. That is k(k^2 + 5) = 6m for some integer m

OpenStudy (anonymous):

Oh i didnt mean our k

OpenStudy (anonymous):

I just mean in general, if something = 6(something else)

OpenStudy (jamesj):

Now you want to show that P(k+1) is true. That is, you want to show that (k+1)( (k+1)^2 + 5 ) is divisible by 6 i.e., (k+1)( (k+1)^2 + 5 ) = 6p for some integer p.

OpenStudy (anonymous):

then something is divisible by 6?

OpenStudy (jamesj):

yes

OpenStudy (jamesj):

If a and b are integers and a = 6b, then a is divisible by 6.

OpenStudy (anonymous):

hm, ok

OpenStudy (jamesj):

I'll leave you in Hero's capable hands or someone else. I've got some other things to do.

OpenStudy (anonymous):

Is hero here?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

we've shown p(1) is true

OpenStudy (anonymous):

Assume \[p(k): (k^{2}+5)k=6m\] is true

OpenStudy (anonymous):

show p(k) => p(k+1)

OpenStudy (anonymous):

\[p(k+1): ((k+1)^{2}+5)(k+1) = 6m\]

OpenStudy (anonymous):

Any tip Zarkon? Im stuck

OpenStudy (zarkon):

multiply out the (k+1)^2

OpenStudy (zarkon):

and stop writing that it is equal to 6m...since that is what we want to show

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[p(k+1): (k ^{2}+2k+6)(k+1)\]

OpenStudy (zarkon):

write is as \[(k ^{2}+5+2k+1)(k+1)\]

OpenStudy (zarkon):

\[[(k ^{2}+5)+(2k+1)](k+1)\]

OpenStudy (zarkon):

now multiply this out ...but keep the \[k^2+5\] as together

OpenStudy (anonymous):

\[(k ^{2}+5)k+(k ^{2}+5)+(2k+1)k+(2k+1)\]

OpenStudy (zarkon):

expand this part ... \[(k ^{2}+5)+(2k+1)k+(2k+1)\]

OpenStudy (anonymous):

\[3k ^{2}+3k+6\]

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

so you have \[(k ^{2}+5)k+3k ^{2}+3k+6\] write as \[(k ^{2}+5)k+3k(k+1)+6\]

OpenStudy (zarkon):

what can you tell me about the above statement

OpenStudy (anonymous):

cubic...

OpenStudy (anonymous):

not enough to say it = 6m, which is what we want right?

OpenStudy (zarkon):

what doe we know about \[(k ^{2}+5)k\]

OpenStudy (anonymous):

=6m

OpenStudy (zarkon):

what about 6

OpenStudy (anonymous):

ah yea

OpenStudy (zarkon):

what about \[3k(k+1)\]

OpenStudy (anonymous):

so we need to show 3k(k+1) = 6m

OpenStudy (zarkon):

which is easy

OpenStudy (anonymous):

3(1)((1)+1) = 3(2) = 6

OpenStudy (zarkon):

\[k(k+1)\] is always even

OpenStudy (anonymous):

So that is sufficient proof

OpenStudy (anonymous):

Ok

OpenStudy (zarkon):

thus \[2|k(k+1)\Rightarrow 6|3k(k+1)\]

OpenStudy (anonymous):

2 Divides? whats that for?

OpenStudy (anonymous):

we only need to show 6 divides

OpenStudy (zarkon):

I'm just pointing out why we get that 6 divides what we want

OpenStudy (anonymous):

So it would be

OpenStudy (anonymous):

Thus \[(k ^{2}+5)k = 6m \rightarrow ((k+1)^{2}+5)(k+1) = 6m\]

OpenStudy (anonymous):

Where the arrow means implies

OpenStudy (zarkon):

I wouldn't use the same letter m

OpenStudy (anonymous):

Im not going to use that at all im just going to say is divisible by 6

OpenStudy (anonymous):

Is that ok to do?

OpenStudy (zarkon):

sure

OpenStudy (anonymous):

Thanks for the help

OpenStudy (zarkon):

np

OpenStudy (anonymous):

Mind checking out a solved question : http://openstudy.com/#/updates/4ee24839e4b0a50f5c5629c8